A projectile is thrown with velocity U = 20 m/s + 5% at an angle 60o. If the projectile falls back on the ground at the same level then which of following can not be a possible answer for range. [Consider g = 10m/s2]


        (a)    39.0 m


        (b)    37.5 m


        (c)    34.6 m


        (d)    32.0 m


Grade:upto college level

1 Answers

Simran Bhatia
10 years ago

(a)

    R = (?20?^2×sin???120?^o ?)/g=20√3=?R/R=24U/U

    =>R = (2×5)/100×200√3=R/R=24U/U

    20√3-2√3<R<20√3+2√3

    => 31.1 m < R < 38.1 m